PEGCOR has a​ P/E ratio of 14.2. Earnings per share are $2.73​,and the expected EPS 5 years from today are $4.19.Calculate the PEG ratio.​ (Refer to Chapter 4 if​necessary.)

The PEG ratio is____
Answer #1

EPS0 = $2.73
EPS5 = $4.19

EPS5 = EPS0 * (1 + Growth Rate)^5
$4.19 = $2.73 * (1 + Growth Rate)^5
1.5348 = (1 + Growth Rate)^5
1.0895 = 1 + Growth Rate
Growth Rate = 0.0895 or 8.95%

PEG Ratio = P/E Ratio / Growth Rate
PEG Ratio = 14.20 / 8.95
PEG Ratio = 1.59